# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ Appendix B Mathematical Tools and Useful Results
$\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}=\frac{1}{M}\sum_{i,j}\mathbb{E}Z_{j,i}^{*}[{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}]_{j,i}=\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{M}-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z)\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}$ (76)
$-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}}{L}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}$
=$\overset{(a)}{=}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{M}-\frac{\delta_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}-\frac{\omega\overline{\omega}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}}{L}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})$
$=(\frac{L}{M}-\omega\overline{\omega})\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{L}-\frac{\delta_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$
where step $(a)$ in (76) follows from (64) in Lemma 3 and the variance control in (74). By substituting (75) into (76) to replace $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}$ and solving $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$, we have
$\frac{1}{M}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}=\frac{(\frac{L}{M}-\omega\overline{\omega})}{1+\delta_{z}\overline{\omega}\overline{\omega}_{z}}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{L}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})=\frac{\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{L}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}).$ (77)
@@
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Q}}(z)}{L}=\frac{\frac{L\delta_{z}^{2}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})}{(\frac{L}{M}-\omega\overline{\omega}+\frac{L\sigma^{2}\delta_{z}}{M}-\frac{\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})})}=\frac{\delta_{z}}{\sigma^{2}+\frac{M\omega\overline{\omega}\overline{\omega}_{z}}{L\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})$ (78)
=$\overset{(a)}{=}\frac{\delta_{z}}{\sigma^{2}+\frac{M\omega_{z}\overline{\omega}\overline{\omega}_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})=\frac{\delta_{z}}{\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$
@@
Next, we turn to evaluate $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}$ in (69). It follows from the resolvent identity
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}=\omega_{z}-\frac{\sigma^{2}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z})=\omega_{z}-\frac{\sigma^{2}\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}\frac{\delta_{z}}{\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z})$ (79)
@@
=$\overset{(b)}{=}\omega_{z}-\frac{\omega_{z}\delta_{z}}{\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}\delta_{z}}{L\delta_{z}^{2}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z})$
$=\frac{\delta\overline{\omega}\omega_{z}\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}\delta_{z}}{L\delta_{z}^{2}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$
where step $(a)$ in (79) follows from the definition of $\Delta_{\sigma^{2}}(z)$ in (71) and step $(b)$ follows from (65). This concludes the proof of (69). Next, we will evaluate $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$ in (70). Notice that $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}$ can be written by integration by parts formula (55) and the variance control as
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}=\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}-\frac{\omega_{z}\overline{\omega}_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$ (80)
$-\frac{\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})$
$=\frac{\overline{\omega}\overline{\omega}_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}).$
@@
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}=\delta\omega_{z}+(1-\frac{M\omega_{z}\overline{\omega}_{z}}{L})\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}-\frac{\delta\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})$ (81)
=$\overset{(a)}{=}\frac{\delta\omega_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}+{\mathcal{O}}(\frac{\mathcal{P}(z^{-1})}{zM^{2}}),$
where $(a)$ is obtained by substituting $\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H}$ using (80) and then solving $\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$. By noticing that $\frac{M\omega(1+\delta\overline{\omega})}{L}=\delta$, we have
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}=\frac{M\omega_{z}\omega(1+\delta\overline{\omega})}{L(1+\delta\overline{\omega}\overline{\omega}_{z})}+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}),$ (82)

### 09 Apr 2024 v5

@@ Appendix B Mathematical Tools and Useful Results
$\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]=\frac{1}{M}\sum_{i,j}\mathbb{E}[Z_{j,i}^{*}[{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}]_{j,i}]=\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{M}-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z))\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})}{M}$ (76)
$-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}})}{L}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}$
=$\overset{(a)}{=}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{M}-\frac{\delta_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})}{M}-\frac{\omega\overline{\omega}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}})]}{L}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}$
$=(\frac{L}{M}-\omega\overline{\omega})\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{L}-\frac{\delta_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$
where step $(a)$ in (76) follows from (64) in Lemma 3 and the variance control in (74). By substituting (75) into (76) to replace $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}$ and solving $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}$, we have
$\frac{1}{M}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]=\frac{(\frac{L}{M}-\omega\overline{\omega})}{1+\delta_{z}\overline{\omega}\overline{\omega}_{z}}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{L}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}=\frac{\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{L}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}.$ (77)
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Q}}(z))]}{L}=\frac{\frac{L\delta_{z}^{2}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})}{(\frac{L}{M}-\omega\overline{\omega}+\frac{L\sigma^{2}\delta_{z}}{M}-\frac{\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})})}=\frac{\delta_{z}}{\sigma^{2}+\frac{M\omega\overline{\omega}\overline{\omega}_{z}}{L\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}$ (78)
=$\overset{(a)}{=}\frac{\delta_{z}}{\sigma^{2}+\frac{M\omega_{z}\overline{\omega}\overline{\omega}_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}=\frac{\delta_{z}}{\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$
@@
Next, we turn to evaluate $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}$ in (69). It follows from the resolvent identity
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}=\omega_{z}-\frac{\sigma^{2}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}\Bigr{)}=\omega_{z}-\frac{\sigma^{2}\omega}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})}\frac{\delta_{z}}{\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}\Bigr{)}$ (79)
@@
=$\overset{(b)}{=}\omega_{z}-\frac{\omega_{z}\delta_{z}}{\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}\delta_{z}}{L\delta_{z}^{2}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}\Bigr{)}$
$=\frac{\delta\overline{\omega}\omega_{z}\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}\delta_{z}}{L\delta_{z}^{2}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$
where step $(a)$ in (79) follows from the definition of $\Delta_{\sigma^{2}}(z)$ in (71) and step $(b)$ follows from (65). This concludes the proof of (69). Next, we will evaluate $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}$ in (70). Notice that $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}$ can be written by integration by parts formula (55) and the variance control as
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}=\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}-\frac{\omega_{z}\overline{\omega}_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}$ (80)
$-\frac{\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}$
$=\frac{\overline{\omega}\overline{\omega}_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}.$
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}=\delta\omega_{z}+(1-\frac{M\omega_{z}\overline{\omega}_{z}}{L})\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}-\frac{\delta\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}$ (81)
=$\overset{(a)}{=}\frac{\delta\omega_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(z^{-1})}{zM^{2}}\Bigr{)},$
where $(a)$ is obtained by substituting $\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{H}}{\mathbb{H}}^{H})]$ using (80) and then solving $\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}$. By noticing that $\frac{M\omega(1+\delta\overline{\omega})}{L}=\delta$, we have
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}=\frac{M\omega_{z}\omega(1+\delta\overline{\omega})}{L(1+\delta\overline{\omega}\overline{\omega}_{z})}+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}),$ (82)